OpenStudy (anonymous):

Two angles with measures (2x^2+3x-5)deg and (x^2+11x-7)deg are supplementary. Find the value of x and then the measure of the each angle

OpenStudy (campbell_st):

well they add to 180 being supplementary so \[2x^2 + 3x -5 + x^2 + 11x - 7 = 180\] collect like terms gives \[3x^2 + 14x - 12 = 180\] or \[3x^2 + 14x - 192 = 0\] just solve the quadratic. General Quadratic formula may be the easiest